# DeStein: Navigating Detoxification of Language Models via Universal Steering Pairs and Head-wise Activation Fusion

### v1

@@ 8 Ethics Statement
In our research, we are careful with the sensitive datasets and the potential for offensive content generation by language models (LMs), aiming to understand and mitigate toxic language dissemination. We recognize the risk that our methodologies for refining LMs to reduce toxic outputs could be misused to create offensive content, a possibility at odds with our ethical commitment to improving LM content safety and integrity. We firmly advocate for the responsible use of LMs, ensuring our work contributes positively to society and opposes any application that promotes harmful communication.
@@ Appendix B Baselines details
To ensure comparability, we rescore the toxicity scores, making certain that they adhere to the same version of the Perspective API (Pozzobon et al., 2023a).
@@
DExperts Liu et al. (2021), a decodingtime method for controlled text generation that combines a pretrained language model with “expert” LMs and/or “anti-expert” LMs in a product of experts. Intuitively, under the ensemble, tokens only get high probability if they are considered likely by the experts and unlikely by the anti-experts.
DisCup, a novel CTG approach that integrates the attribute knowledge of a discriminator to optimize control-prompts, guiding a frozen CLM in producing attribute-specific texts. Initially, the frozen CLM model, known for its ability to generate diverse texts, is tasked with generating next-token candidates based on context to ensure token prediction diversity. Subsequently, an attribute-discriminator is employed to select desired or undesired tokens from these candidates, incorporating inter-attribute knowledge. Finally, these traits are unified through an unlikelihood objective for prompt-tuning.
GOODTRIEVER, based on KNN-LM, integrates a retrieval-based approach at decoding, enabling it to facilitate toxicity-controlled text generation. The retrieval corpus is constructed based on the Jigsaw Unintended Bias dataset.
Self-Detoxify builds upon this concept by crafting a technique to discern the direction of toxification from the typical generation process to the altered one induced by the negative prefix. Subsequently, they guide the generation process towards the opposite direction by controlling how information flows within the attention layers. In the experiments showcased by this method, we utilize Liu et al. (2021) released checkpoint as the base model. To ensure a fair comparison, we employ two distinct setups: one based on the checkpoints and another based on the original GPT2-large model. The parameters are set accordingly to yield the best results observed in the experiments: $\alpha=2$, $\beta=2.5$ for GPT2-large. To enhance the performance of the large model, we made minor adjustments to its parameters: $\alpha=2$, $\beta=1.5$.
@@ Appendix C The details of prompts
Please choose the better one. If the quality of two texts is the same, please output ’D’. Only ’A’, ’B’, or ’D’ is allowed. Please make an effort to discern, and strive to avoid outputting ’D’ as much as possible.
@@ Appendix D Experiments details
Unconditional generation details. For GPT2-large, we use nucleus sampling with p=0.9, a generation temperature of 1.0, and a maximum generation length of 100. For LLMs, due to the reduced likelihood of toxic outputs from large models, generating 10k texts unconditionally fails to yield a sufficient amount of toxic sentences. Moreover, our focus is primarily on demonstrating the scalability of our approach on large models rather than meticulously designing experiments to achieve optimal results. Therefore, to conserve computational resources, we employ some toxic-inducing techniques for generation. Specifically, we randomly select 1000 toxic sentences from the ParaDetox dataset(Logacheva et al., 2022) as inducing prompts and use the same generation parameters as mentioned above for toxic text generation.
@@
Implementation details. Due to the fact that sampling methods can affect PPL and diversity, for a fairer comparison, the parameters of our nucleus sampling are kept consistent across all methods. We use the hyperparameters of top-k=0, top-p=0.9 and temperature=1.0.
Appendix E Ablation experiments
Based on our findings that the number of prompts did not significantly affect the results, we extracted a balanced test dataset with 50 toxic and 50 non-toxic prompts for evaluation.
@@ Appendix E Ablation experiments
Ablation experiments on attention heads. Specifically, we considered the following variants of our method: 1) Setting $\alpha_{prob}$ to a fixed value of 0.5; 2) Sorting $\alpha_{prob}$ by magnitude and retaining only the top half; 3) Contrary to the previous variant, keeping only the bottom half.
@@
Table 6: The experiment involves the dissolution of attention heads during the head-wise fusion.
Model Toxicity ($\downarrow$) Fluency ($\downarrow$) Diversity ($\uparrow$)
Avg. Max. Tox. Tox. Prob. PPL Dist-1 Dist-2 Dist-3
OURs (bottom) 0.315 0.16 31.032 0.577 0.859 0.858
OURs (fixed=0.5) 0.207 0.04 39.434 0.569 0.859 0.860
OURs (top) 0.262 0.10 33.163 0.577 0.858 0.859
The results, as shown in Table 6, validate the intuition behind Probing-based Weights, suggesting that their application to fusion is logical. Indeed, the probing technique is capable of identifying activation spaces that are more effective for detoxifying language models.
Ablation experiment on data. To demonstrate the contribution of my self-supervised parallel data to toxicity removal, we conducted the following experiments: 1) We performed toxicity removal using non-parallel data, ensuring fairness by using toxic sentences identical to those in the parallel dataset; 2) We conducted toxicity removal using non-self-induced data, with parallel pairs sourced from ParaDetox. Note that apart from the differences in data, all other experimental settings remained consistent. The results are shown in Table 7.
Table 7: Ablation experiments on data.
Model Toxicity ($\downarrow$) Fluency ($\downarrow$) Diversity ($\uparrow$)
Avg. Max. Tox. Tox. Prob. PPL Dist-1 Dist-2 Dist-3
OURs (non-self-induced) 0.327 0.19 32.145 0.566 0.862 0.863
OURs (non-parallel) 0.216 0.07 41.567 0.564 0.855 0.863
OURs 0.203 0.04 39.405 0.569 0.858 0.860
Appendix F Additional analysis of toxic and non-toxic prompts
@@
Consistent with the general definition, we define prompts with toxicity scores greater than 0.5 from the Perspective API as toxic prompts, and those with scores equal to or less than 0.5 as non-toxic prompts. We experimented with 5k toxic prompts and 5k non-toxic prompts. We observed that our method outperformed all previous methods in terms of PPL for toxic prompts. Furthermore, it is reasonable for the PPL of results corresponding to toxic prompts to increase because detoxification can lead to steering of language style in both continuation and prompt. However, we observed an increase in PPL for results corresponding to non-toxic prompts as well, with some increases even surpassing those of toxic prompts, which is not desirable. This phenomenon is not exclusive to our method; nearly all methods exhibit this behavior, except for GeDi.
We analyze the reasons behind this phenomenon for our method. This discrepancy arises because our method treats toxic and non-toxic prompts indiscriminately. During inference, steering using detox vectors is reasonable if the activation space’s vector distribution falls within the toxic region, explaining our method’s effective detoxification. However, applying the same detox operation in regions where vectors predominantly represent non-toxic prompts leads to increased PPL. Consequently, we suggest integrating our method with toxicity classifiers to achieve better results. Moreover, our method introduces minimal additional parameters, which are easily integrable, making it suitable for practical deployment alongside pre-existing toxicity classifiers. This integration represents a viable security measure for large model deployment, which is a direction for our future work.
Table 8: Automatic evaluation results on toxic and nontoxic prompts with GPT2-large. Best models are shown in bold and 2nd best models are underlined.
@@
Toxicity ($\downarrow$) Fluency ($\downarrow$) Toxicity ($\downarrow$) Fluency ($\downarrow$)
Exp. Max. Tox. Tox. Prob. PPL Exp. Max. Tox. Tox. Prob. PPL
@@
OURs 0.264 0.111 41.002 0.142 0.012 34.615
Table 9: Automatic evaluation results on toxic and nontoxic prompts with LLMs. Best models are shown in bold.
@@
Toxicity ($\downarrow$) Fluency ($\downarrow$) Toxicity ($\downarrow$) Fluency ($\downarrow$)
Exp. Max. Tox. Tox. Prob. PPL Exp. Max. Tox. Tox. Prob. PPL
LLaMA-2-7b 0.696 0.833 18.690 0.382 0.267 16.684
@@
OURs 0.354 0.199 93.306 0.360 0.219 12.900
Appendix G Case study
@@
In this section, we provide some specific examples for different detoxification methods. In our experimental setup, each prompt corresponds to generating 25 continuations, from which we select the most representative ones among the top three in terms of toxicity. The examples could be seen in Table 10.
Table 10: Example sentences that are generated by each model.
@@
OURs “and will be on your tail. If they lose interest in discussing their individual interest, then a good” 0.158
@@
LLaMA-2-7b BASE “we see you. We are going to kill you soon. What happened to the picture of the” 0.782
@@
OURs “we meet. There will be a death sentence in the form of capital punishment, or” 0.323
@@
OURs “you’re invited out of phase. Well, what they do *will* annoy the hell out” 0.373
@@
OURs “be available to the channel. Because it may be another competitor, trying to steal that view from you” 0.238

### v3

@@ 8 Ethics Statement
In our experiments, we are careful with the sensitive datasets and the potential for offensive content generation by language models (LMs), aiming to understand and mitigate toxic language dissemination. We recognize the risk that our method for refining LMs to reduce toxic outputs could be misused to create offensive content, a possibility at odds with our ethical commitment to improving LM content safety and integrity. We firmly advocate for the responsible use of LMs, ensuring our work contributes positively to society and opposes any application that promotes harmful communication.
@@ Appendix B Baselines details
For comparability, we have rescored the toxicity scores to ensure they adhere to the same version of the Perspective API (Pozzobon et al., 2023a).
@@
DEXPERTS (Liu et al., 2021), a decoding-based method for controlled text generation that combines a pre-trained language model with “expert” LMs and “anti-expert” LMs in a product of experts. Intuitively, under the ensemble, tokens only get a high probability if they are considered likely by the experts and unlikely by the anti-experts.
DisCup, a novel CTG approach that integrates the attribute knowledge of a discriminator to optimize control prompts, guiding a frozen CLM in producing attribute-specific texts. Initially, the frozen CLM model, known for its ability to generate diverse texts, is tasked with generating next-token candidates based on context to ensure token prediction diversity. Subsequently, an attribute-discriminator is employed to select desired or undesired tokens from these candidates, incorporating inter-attribute knowledge. Finally, these traits are unified through an unlikelihood objective for prompt-tuning.
GOODTRIEVER, based on KNN-LM, integrates a retrieval-based approach to decoding, enabling it to facilitate toxicity-controlled text generation. The retrieval corpus is constructed based on the Jigsaw Unintended Bias dataset.
Self-Detoxify builds upon this concept by crafting a technique to discern the direction of toxification from the typical generation process to the altered one induced by the negative prefix. Subsequently, they guide the generation process in the opposite direction by controlling how information flows within the attention layers. In the experiments showcased by this method, we utilize Liu et al. (2021) released checkpoint as the base model. For a fair comparison, we employ two distinct setups: one based on the checkpoints and another based on the original GPT2-large model. The parameters are set accordingly to yield the best results observed in the experiments: $\alpha=2$, $\beta=2.5$ for GPT2-large. To enhance the performance of the large model, we made minor adjustments to its parameters: $\alpha=2$, $\beta=1.5$.
@@ Appendix C The details of prompts
Please choose the better one. If the quality of the two texts is the same, please output ’D’. Only ’A’, ’B’, or ’D’ are allowed. Please make an effort to discern, and strive to avoid outputting ’D’ as much as possible.
@@ Appendix D Experiments details
Unconditional generation details. For GPT2-large, we use nucleus sampling with p=0.9, a generation temperature of 1.0, and a maximum generation length of 100. For LLMs, due to the reduced likelihood of toxic outputs from large models, generating 10k texts unconditionally fails to yield a sufficient amount of toxic samples. Moreover, our focus is primarily on demonstrating the scalability of our approach on large models rather than meticulously designing experiments to achieve optimal results. Therefore, to conserve computational resources, we employ some toxic-inducing techniques for generation. Specifically, we randomly select 1000 toxic samples from the ParaDetox dataset (Logacheva et al., 2022) as inducing prompts and use the same generation parameters as mentioned above for toxic text generation.
@@
Implementation details. Because sampling methods can affect PPL and diversity, the parameters of our nucleus sampling are kept consistent across all methods for fair comparison. We use the hyperparameters of top-k=0, top-p=0.9, and temperature=1.0.
Appendix E Parameter count calculation
In this section, we describe the method for calculating memory usage in our experiments. Specifically, the calculation process is Total Memory (TM) $=N_{l}\times N_{h}\times D_{h}\times B$, where $N_{l}$ is the number of layers, $N_{h}$ is the number of attention heads per layer, $D_{h}$ is the output vector dimension per head, and $B$ is bytes per value (4 for float32). The results of our memory usage calculations are summarized in Table 8. Notably, the additional memory usage incurred by our method is minimal, even for LLMs.
Model $N_{l}$ $N_{h}$ $D_{h}$ memory (single head) memory(all)
GPT2-large 36 20 64 256 bytes 180 KB
GPT2-XL(1.3B) 48 25 64 256 bytes 300 KB
LLaMA2-7B(OPT-6.7B and MPT-7B) 32 32 128 512 bytes 512 KB
LLaMA2-13B 40 40 128 512 bytes 800 KB
Table 8: Memory usage for various models.
Appendix F Additional analysis of ablation study
The quality of parallel steering pairs. We examine the quality of parallel steering pairs ($m$) and their effect on detoxification. The results are shown in Table 9. Our results demonstrate that constructing and filtering data according to our method yields significantly improved performance with as few as 20 randomly selected pairs. Further increasing the number of pairs did not yield substantial improvements. To conserve computational and API resources, we opted to use 20 pairs for GPT2-large. Additionally, repeated random selections, regardless of the number of pairs chosen, consistently showed good performance, indirectly highlighting the robustness of our approach.
Value Toxicity $\downarrow$ Fluency $\downarrow$ Diversity $\uparrow$
EMT TP PPL Dist-1 Dist-2 Dist-3
$m$=5 0.307 0.14 38.222 0.577 0.858 0.858
$m$=10 0.209 0.05 51.869 0.562 0.849 0.857
$m$=20 0.203 0.04 39.405 0.569 0.858 0.860
$m$=40 0.229 0.08 43.088 0.585 0.862 0.862
$m$=60 0.213 0.08 43.364 0.588 0.862 0.862
Table 9: Evaluation results of the experiments conducted with varying $m$ on GPT2-large.
The detoxification strength ($\alpha_{contr}$). The detailed impact of the toxicity control parameter $\alpha_{contr}$ on performance is shown in Table 10. The results indicate that our approach can achieve flexible toxicity control.
Value Toxicity $\downarrow$ Fluency $\downarrow$ Diversity $\uparrow$
EMT TP PPL Dist-1 Dist-2 Dist-3
$\alpha_{contr}$=0.1 0.426 0.33 26.972 0.584 0.857 0.854
$\alpha_{contr}$=0.3 0.270 0.11 32.113 0.576 0.857 0.857
$\alpha_{contr}$=0.4 0.203 0.04 39.405 0.569 0.858 0.860
$\alpha_{contr}$=0.6 0.107 0.01 66.363 0.557 0.859 0.864
Table 10: Evaluation results of the experiments conducted with varying $\alpha_{contr}$ on GPT2-large.
@@ Appendix F Additional analysis of ablation study
Head-wise coefficients. Specifically, we considered the following variants of our method: 1) Sorting $\alpha_{prob}$ by magnitude and retaining only the top half; 2) Contrary to the previous variant, keeping only the bottom half. The results, as shown in Table 11, validate the intuition behind probing-based Weights, suggesting that their fusion application is logical.
Model Toxicity $\downarrow$ Fluency $\downarrow$ Diversity $\uparrow$
EMT TP PPL Dist-1 Dist-2 Dist-3
DeStein(bottom) 0.315 0.16 31.032 0.577 0.859 0.858
DeStein(top) 0.262 0.10 33.163 0.577 0.858 0.859
@@
Table 11: Evaluation results of the dissolution of attention heads during head-wise fusion with GPT2-large.
Appendix G Additional analysis of toxic and non-toxic prompts
@@
Consistent with the general definition, we define prompts with toxicity scores greater than 0.5 from the Perspective API as toxic prompts and those with scores equal to or less than 0.5 as non-toxic prompts. We experimented with 5k toxic prompts and 5k non-toxic prompts. We observed that our method outperformed previous methods in terms of PPL for toxic prompts. Furthermore, it is reasonable for the PPL of results corresponding to toxic prompts to increase because detoxification can lead to steering of language style in both continuation and prompt. However, we observed an increase in PPL for results corresponding to non-toxic prompts as well, with some increases even surpassing those of toxic prompts, which is not desirable. This phenomenon is not exclusive to our method; nearly all methods exhibit this behavior, except for GeDi.
We analyze the reasons behind this phenomenon for our method. This discrepancy arises because our method treats toxic and non-toxic prompts indiscriminately. During inference, steering using the detoxification vector is reasonable if the activation space’s vector distribution falls within the toxic region, explaining our method’s effective detoxification. However, applying the same detoxification operation in regions where vectors predominantly represent non-toxic prompts leads to increased PPL. Consequently, we suggest integrating our method with toxicity classifiers to achieve better results. Moreover, our method introduces minimal additional parameters, which are easily integrable, making it suitable for practical deployment alongside pre-existing toxicity classifiers. This integration represents a viable security measure for large model deployment, which is a direction for our future work.
@@
Toxicity $\downarrow$ Fluency $\downarrow$ Toxicity $\downarrow$ Fluency $\downarrow$
EMT TP PPL EMT TP PPL
@@
DeStein 0.264 0.111 41.002 0.142 0.012 34.615
Table 12: Evaluation results on toxic and nontoxic prompts with GPT2-large. The best results are shown in bold, and the 2nd best results are underlined.
@@
Toxicity $\downarrow$ Fluency $\downarrow$ Toxicity $\downarrow$ Fluency $\downarrow$
EMT TP PPL EMT TP PPL
LLaMA2-7b 0.696 0.833 18.690 0.382 0.267 16.684
@@
DeStein 0.393 0.292 19.185 0.190 0.022 16.281
Table 13: Evaluation results on toxic and nontoxic prompts with LLMs. The best results are shown in bold.
Appendix H Case study
@@
In this section, we provide some specific examples of different detoxification methods. In our experimental setup, each prompt corresponds to generating 25 continuations, from which we select the most representative ones among the top three in terms of toxicity. The examples can be seen in Table 14.
@@
DeStein “and will be on your tail. If they lose interest in discussing their individual interest, then a good” 0.158
@@
LLaMA2-7b Base “we see you. We are going to kill you soon. What happened to the picture of the” 0.782
@@
DeStein “we meet. There will be a death sentence in the form of capital punishment, or” 0.323
@@
DeStein “you’re invited out of phase. Well, what they do *will* annoy the hell out” 0.373
@@
DeStein “be available to the channel. Because it may be another competitor, trying to steal that view from you” 0.238
Table 14: Example sentences generated by each model.